# Explainable Data-driven Modeling of Adsorption Energy in Heterogeneous Catalysis

### v1

@@ Abstract
The increasing popularity of machine learning (ML) in catalysis has spurred interest in leveraging these techniques to enhance catalyst design. Despite the promise of machine learning approaches, they often function as black boxes, leaving a gap between physics-based studies and data-driven methodologies. Our study aims to bridge this gap by integrating ML techniques with eXplainable AI (XAI), specifically employing two techniques: post-hoc XAI analysis and Symbolic Regression, to unravel the correlation between adsorption energy and the properties of the adsorbate-catalyst system. Leveraging a large dataset such as the Open Catalyst Dataset (OC20), we employ a combination of shallow ML techniques and XAI methodologies. Our investigation involves utilizing multiple shallow machine learning techniques to predict adsorption energy followed by Post-hoc analysis for feature importance, inter-feature correlations, and the influence of various features values on prediction of adsorption energy. The Post-hoc analysis reveals that adsorbate properties exert a greater influence than catalyst properties in our dataset. The top five features with the most importance, based on higher Shapley values, are adsorbate electronegativity, the number of adsorbate atoms, catalyst electronegativity, effective coordination number, and the sum of atomic numbers of the adsorbate molecule. There is a positive correlation between catalyst and adsorbate electronegativity with the prediction of adsorption energy. Additionally, Symbolic Regression, which helps elucidate mathematical expressions describing the relationship between input features and adsorption energy, yields results consistent with SHAP analysis. It deduces a mathematical relationship indicating that the square of the catalyst electronegativity is directly proportional to the adsorption energy. These consistent correlations resemble those derived from physics-based equations in previous research. Our work establishes a robust framework that integrates ML techniques with XAI, leveraging large datasets like OC20 to enhance catalyst design through model explainability.
@@ 1 Introduction
Catalysis plays a crucial role in chemical reactions and finds wide applications in industries, driving over 90% of manufacturing processes1. These processes consume about 25% of industrial energy, with heterogeneous catalysis being particularly energy-intensive. 2. To address environmental concerns and promote sustainability, it is essential to enhance the efficiency of the reaction through optimal catalyst design, which aims for higher selectivity and activity.
@@
In catalyst design and development, a critical focus lies on understanding adsorption energy due to its strong correlation with catalytic reactivity, governing the interaction between reactants and catalysts 3. The Brønsted-Evans-Polanyi (BEP) relations illustrate volcano-shaped relationships between catalytic activity and adsorption energies, highlighting it as a pivotal descriptor 4, 5, 6, 7. The BEP relation demonstrates a linear relationship between activation barrier, $E_{a}$, and adsorption energy, $\Delta E_{\text{ads}}$, in heterogeneous catalysis, as depicted in Equation 1. This linear relationship implies that the catalyst’s reactivity can be primarily characterized by a descriptor, $\Delta E_{\text{ads}}$. Modeling adsorption energy provides insights into reactivity and aids in catalyst screening, thus understanding the factors affecting adsorption energy is essential to identify the optimal catalyst for a target reaction.
$E_{a}=\gamma\Delta E_{\text{ads}}+\xi$ (1)
@@
Density Functional Theory (DFT) is a well-established quantum chemistry-based method widely used to calculate the properties of atomic systems, including adsorption energy in catalysis. Despite its high accuracy, DFT requires substantial computational resources. Consequently, evaluating a large number of materials for catalyst design using expensive quantum chemistry-based approaches like DFT for high-throughput screening becomes infeasible. As a solution, researchers are turning to machine learning approaches, which offer the potential to accurately predict properties at a reduced computational cost. Notably, extensive databases such as the Open Catalyst 2020 (OC20) dataset 8 comprise over 1.2 million DFT relaxations of adsorbate-catalyst systems (approximately 250 million single-point calculations) across a substantially broader structure and chemistry space than previously realized. These datasets serve as the foundation for training and deploying machine learning models, aimed at predicting the energy and interatomic forces of the adsorbate-catalyst systems. The extensive availability of material databases like OC20 for adsorbate-catalyst systems has significantly accelerated the adoption of ML approaches in catalysis, particularly Graph Neural Networks (GNNs) 9, 10, 11, 12, 13, 14, 15.
@@
Despite the rise of ML approaches in atomic modeling, their black-box nature obscures the underlying physical insights related to energy and force predictions. This lack of interpretability limits the further application of ML methods, as the contributing factors to predictions remain unclear. Additionally, since ML predictions inherently come with a certain level of uncertainty, it is more beneficial to gain insights about catalyst systems from ML modeling rather than solely relying on prediction values 16. To address this challenge, this paper introduces the use of eXplainable Artificial Intelligence (XAI), which uncovers the underlying correlations behind predictions and provides human-understandable explanations, thereby enhancing domain-specific knowledge17. To achieve explainability, we implement two strategies: firstly, training shallow machine learning models such as Adaboost Regression, XGBoost Regression, Support Vector Regression (SVR), Kernel Ridge Regression (KRR), and Least Absolute Shrinkage and Selection Operator (LASSO) Regression accompanied by post-hoc XAI analysis; secondly, applying Symbolic Regression to gain better understanding of relation between input features and adsorption energy. Both approaches use dataset sourced from the OC20 dataset. The former strategy highlights the importance and correlations of features, while the latter generates mathematical equations that directly expose the potential relationships between input features and target labels.
@@ 2.1 Framework
Our study leverages XAI to extract meaningful insights about the relationship between the characteristics of adsorbate-catalyst systems and the adsorption energy, which is the label value for ML models. We utilize XAI in two distinct ways as seen in Figure 1. Firstly, through shallow machine learning techniques alongside post-hoc XAI analysis. The post-hoc XAI analysis involves using the SHAP library 18. SHAP analysis provides the importance of individual features based on Shapley values. Secondly, we employ Symbolic Regression (SR) to derive mathematical equations to compute the adsorption energy based on the given input features. These equations are compared to equations obtained from traditional theory or experimental-based methods, highlighting their potential complementarity. The feature importance and correlation among features from SHAP analysis, along with the mathematical equations obtained from the Symbolic regression technique, provide better insights into the relation of adsorption energy with structural properties.
The adsorption energies and feature values like electronegativity, coordination number, and so forth, for both methods are extracted from the adsorbate-catalyst systems in the Open Catalyst 2020 (OC20) dataset. The details of the feature extraction are elaborated in the following section. Integrating ML techniques with XAI using large datasets like the OCP dataset lays the groundwork for effective catalyst design and opens promising avenues for future research.
@@ 2.2 Selection of Features
The features are carefully chosen from previous physics-based modeling research regarding adsorption energy computation. The features are categorized into two groups: adsorbate-specific and catalytic-specific features, as seen in Table 1. Each feature provides insights into electronic structure, bonding characteristics, and energetic properties of adsorption sites and catalysts. For instance, local electronegativity is determined as the geometric mean of the Pauling electronegativity of metal atoms within the first neighboring shell, offering insight into the local reactivity of catalyst surfaces19 as follows:
@@
where $\chi_{j}^{0}$ is the electronegativity of atom j and N is the total number of atoms within the first neighboring shell including the adsorption site i. While, site type represented numerical classifications of different adsorption site configurations such as bridge sites, atop sites, hollow-HPC sites, etc.
@@
Number atoms in adsorbate $N$ Amos et al.21 Total number of atoms present in adsorbate molecule
Sum of adsorbate atomic numbers $Z_{ads}$ Tran et al.22 Sum of atomic numbers of all atoms in adsorbate molecule
Center atom coordination number $CN_{ads}$ Calle-Vallejo et al.23 Coordination number of the central atom in adsorbate molecule
Catalyst Catalyst electronegativity $\chi_{cat}$ Ishioka et al.20 Average electronegativity of the interacting surface atoms of catalyst
Effective coordination number $CN_{cat}$ Calle-Vallejo et al.23 Average coordination number of the interacting surface atoms of catalyst
Site type $S$ Roy et al. 24 Type of site where the adsorbate interacts with catalyst surface
@@
Miller index $hkl$ Chen et al.27 Miller index of catalyst
Formation energy $\Delta H_{f}$ Ha et al.28 Formation energy of catalyst
@@ 2.3 Post-hoc XAI analysis
The dataset, comprising 2886 adsorbate-catalyst systems, is divided into an 80:20 ratio for training and testing purposes, with input features and the adsorption energy label provided to the machine learning models.Details about the dataset and data selection are elaborated in the Methods section. The performance outcomes of the models are presented in Table 2. The outcomes demonstrated by all models exhibit comparable performance with MAE values ranging from 0.446 to 0.617. Notably, Adaboost regression stands out as the best performer, demonstrating the lowest MAE value. Hence, it is used as the reference model for SHAP analysis during the post-hoc analysis.
Table 2: Performance of different regression models
@@
In the post-hoc analysis, the primary aspect of model explainability is carried out through SHAP analysis 31. The SHAP library has been widely used in various domains, including materials science, to interpret black box machine learning models 18. This technique computes Shapley values for all features, thereby elucidating each feature’s contribution to the prediction of adsorption energy by assessing the impact of each feature across different combinations of feature values. Further details on SHAP approach is provided in the Methods section. Visualizations such as the summary plot and radar chart, as depicted in Figure 3(a) and Figure 3(b), aid in determining feature importance, while the beeswarm plot in Figure 3(c), offers insights into individual feature value impacts on adsorption energy. Additionally, scatter plots are employed to reveal feature-feature correlations in Figure 4. Moreover, a correlation matrix is constructed in the process to further understand feature interactions, as shown in Figure 2(b).
@@
Figure 2: Performance Evaluation and Feature Correlation: a. parity plot illustrating the performance of Adaboost Regression with Base estimator as Random Forest Regressor (Best model). MAE values are calculated for systems with O and H group and C1 group. b. Correlation matrix depicting the relationships between input features. High values of correlation coefficient have been found between Local electronegativity and Catalyst electronegativity, Site type and coordination number of adsorbate molecule.
From the radar chart and summary plot in Figure 3, we can deduce that the top 5 features with the most importance (higher Shapley values) are adsorbate electronegativity, number of adsorbate atoms, catalyst electronegativity, effective coordination number, and the sum of atomic numbers of adsorbate molecule, with Shapley values of 0.42, 0.39, 0.27, 0.15, and 0.08, respectively. Furthermore, it is evident that adsorbate properties exert a greater influence than catalyst properties. This observation aligns with findings from the language model-based predictive model32, which is specifically designed to predict the adsorption energies of adsorbate-catalyst systems. The attention score from the language model approach also shows that the model focuses more on adsorbate-related tokens. Likewise, our SHAP analysis demonstrates that the adsorption energy predictions rely more on the adsorbate-related features.
@@
Previous studies have shown varying relationships between catalyst electronegativity and adsorption energy. Gao, Wang, et al. proposed an inverse relationship specifically for transition metal catalysts, where higher electronegativity leads to lower adsorption energy 33. Conversely, Trasatti’s research suggests a direct proportionality between metal catalyst electronegativity and adsorption energy 34.
Our analysis reveals a positive correlation between catalyst electronegativity and adsorption energy. This result is evident in the beeswarm plot analysis shown in Figure 3(c), where significant clustering of data points and feature values for catalyst electronegativity can be observed for positive Shapley values. While Gao and Wang found an inverse relationship in transition metals, our results suggest the opposite, aligning instead with Trasatti’s findings.
It is important to note that our dataset includes not only hydrogen but also oxygen and carbon-based adsorbates with metal catalysts, which may account for the differences observed in comparison to the work of Gao and Wang. This underscores the intricate nature of catalyst-adsorbate interactions, where various factors such as the composition of the dataset and the choice of adsorbates and catalysts play crucial roles in determining the observed outcomes. Additionally, a similar trend is observed for adsorbate electronegativity and local electronegativity, supporting our findings and emphasizing the complexity of catalyst-adsorbate interactions.
@@
The electronegativity of the catalyst bulk can effectively serve as a proxy for surface electronegativity in adsorption studies. This deduction is supported by Figure 4, which explores feature-feature correlations and reveals two notable findings. Firstly, Figure 4(a) demonstrates a positive correlation between mean catalyst electronegativity and local electronegativity. This relationship reflects the influence of catalyst atoms within the first neighboring shell at the adsorption site. It logically follows that the electronegativity of a subset of surface atoms contributes to the overall mean electronegativity of the catalyst lattice. Additionally, a similar correlation is observed in the correlation matrix depicted in Figure 2(b), where local electronegativity and catalyst electronegativity exhibit a high correlation coefficient. This suggests a strong association between surface and bulk electronegativity in our dataset. Such findings are crucial for practical applications since adsorption primarily occurs at the surface. Despite challenges in accurately characterizing surface properties in real-world scenarios, these results indicate that the electronegativity of the catalyst bulk can effectively represent surface electronegativity in adsorption studies, thus simplifying materials research efforts.
The increase in the number of adsorbate atoms correlates with a decrease in overall adsorbate electronegativity, as depicted in Figure 4(b). This trend is evident from the transition in color scheme from orange (indicating high adsorbate electronegativity) to purple (indicating lower adsorbate electronegativity). The decrease in electronegativity could be attributed to the formation of chemical bonds between the adsorbate atoms. As more adsorbate atoms come into proximity, bonding interactions occur, causing a redistribution of electron density within the adsorbate system and thereby lowering the overall electronegativity 35.

### v2

@@ Abstract
The increasing popularity of machine learning (ML) in catalysis has spurred interest in leveraging data-driven approaches to enhance catalyst design. Despite the promise of machine learning approaches, they often function as black boxes, leaving a gap between physics-based studies and data-driven methodologies. Our study aims to bridge this gap by integrating ML techniques with eXplainable AI (XAI), specifically employing two techniques: post-hoc XAI analysis and Symbolic Regression, to unravel the correlation between adsorption energy and the properties of the adsorbate-catalyst system. Leveraging a large dataset such as the Open Catalyst 2020 (OC20) Dataset, we employ a combination of shallow ML techniques and XAI methodologies. Our investigation involves utilizing multiple shallow machine learning techniques to predict adsorption energy followed by post-hoc analysis for feature importance, inter-feature correlations, and the influence of various feature values on prediction of adsorption energy. Features such as adsorbate electronegativity, coordination number of the adsorbing atom, electronegativity of the adsorbing atom, catalyst electronegativity, and the number of atoms in the adsorbate stand out in our post-hoc analysis within our feature set for predicting adsorption energy. There is a positive correlation between catalyst and adsorbate electronegativity with the prediction of adsorption energy. Additionally, Symbolic Regression, which helps elucidate mathematical expressions describing the relationship between these important features and adsorption energy, yields results consistent with physics-based equations in previous research. It deduces a mathematical relationship indicating that coordination number of the adsorbing atom is directly proportional to the adsorption energy. Our work establishes a robust framework that integrates ML techniques with XAI, leveraging large datasets like OC20 to enhance catalyst design through model explainability.
@@ 1 Introduction
Rational design of optimal catalysts remains one of the critical challenges in heterogeneous catalysis 1. While the primary goal is to enhance catalytic performance, the underlying physics spans across nanoscale phenomena, interfacial chemistry, and reaction engineering 2. The multifaceted nature of catalyst surfaces and interfaces makes it particularly difficult to establish universal theoretical frameworks that can reliably predict catalyst reactivity through simple descriptors. Despite these challenges, significant progress has been made in identifying key parameters that govern catalytic behavior.
@@
In catalyst design and development, a critical focus lies on understanding adsorption energy due to its strong correlation with catalytic reactivity, governing the interaction between reactants and catalysts 3. The Brønsted-Evans-Polanyi (BEP) relations illustrate volcano-shaped linear relationships between catalytic activity and adsorption energies, highlighting it as a pivotal descriptor 4, 5, 6, 7. The BEP relation demonstrates a linear relationship between activation barrier, $E_{a}$, and adsorption energy, $E_{\text{ads}}$, in heterogeneous catalysis. This linear relationship implies that the catalyst’s reactivity can be primarily characterized by a descriptor, $E_{\text{ads}}$. Modeling adsorption energy provides insights into reactivity and aids in catalyst screening, thus understanding the factors affecting adsorption energy is essential to identify the optimal catalyst for a target reaction.
@@
Density Functional Theory (DFT) is a quantum chemistry-based method widely used to calculate the properties of atomic systems, including adsorption energy in catalysis. Despite its high accuracy, DFT requires substantial computational resources. Consequently, evaluating a large number of materials for catalyst design using expensive quantum chemistry-based approaches like DFT for high-throughput screening becomes infeasible. As a solution, machine learning approaches are gaining much attention due to their potential to accurately predict properties, such as energy and forces, at a reduced computational cost. Notably, extensive databases such as the OC20 dataset 8 comprise over 1.2 million DFT relaxations of adsorbate-catalyst systems (approximately 250 million single-point calculations) across a substantially broader structure and chemistry space than previously realized. These datasets serve as the foundation for training and deploying machine learning models, aimed at predicting the energy and interatomic forces of the adsorbate-catalyst systems. The extensive availability of material databases like OC20 for adsorbate-catalyst systems has significantly accelerated the adoption of ML approaches in catalysis, particularly Graph Neural Networks (GNNs) 9, 10, 11, 12, 13, 14, 15.
@@
Despite the rise of ML approaches in atomic modeling, their black-box nature obscures the underlying physical insights related to energy and force predictions. This lack of interpretability limits the further application of ML methods, as the contributing factors to predictions remain unclear. Additionally, since ML predictions inherently come with a certain level of uncertainty16, it is more beneficial to gain insights about catalyst systems from ML modeling rather than solely relying on prediction values. To address this challenge, this paper introduces the use of XAI, which uncovers the underlying correlations behind predictions and provides human-understandable explanations, thereby enhancing domain-specific knowledge17. To achieve explainability, we implement two strategies: firstly, training shallow machine learning models such as Adaboost Regression, XGBoost Regression, Support Vector Regression (SVR), Kernel Ridge Regression (KRR), and Least Absolute Shrinkage and Selection Operator (LASSO) Regression accompanied by post-hoc XAI analysis; secondly, applying Symbolic Regression to gain better understanding of relation between top 5 input features and adsorption energy. The former strategy highlights the importance and correlations of features, while the latter generates mathematical equations that directly expose the potential relationships between input features and target label.
@@ 2.1 Framework
Our study leverages XAI to extract meaningful insights about the relationship between the characteristics of adsorbate-catalyst systems and the adsorption energy, which is the label value for ML predictions. We utilize XAI in two distinct ways as seen in Figure 1. First, through shallow machine learning techniques combined with post-hoc XAI analysis. The post-hoc XAI analysis is conducted using the SHAP (SHapley Additive exPlanations) library 18, which provides feature importance based on Shapley values. This feature importance is crucial for understanding which characteristics most strongly influence adsorption energy, helping to guide catalyst design by highlighting key factors that impact system performance. Second, we employ Symbolic Regression to derive mathematical equations to compute the adsorption energy based on the top 5 input features with higher Shapely values. Obtaining mathematical equations to calculate adsorption energy is crucial as they enable a clear, quantitative understanding of complex relationships in data, allowing for efficient predictions, easier integration into simulations, and insights into the governing principles of adsorption phenomena. These equations are compared to equations obtained from traditional theory or experimental-based methods, highlighting their potential complementarity.
The feature importance and correlation among features from SHAP analysis, along with the mathematical equations obtained from the Symbolic regression technique, provide better insights into the relation of adsorption energy with structural properties in a data-driven manner. The adsorption energies and feature values like electronegativity, coordination number, and so forth, for both methods are extracted from the adsorbate-catalyst systems in the OC20 dataset. The details of the feature extraction are elaborated in the following section. Integrating ML with XAI on large datasets, such as the OC20 dataset, yields valuable modeling insights that enhance our understanding of catalyst behavior. These insights can be validated through physics-based studies, providing a robust framework for informed catalyst design and accelerating advancements in the field.
@@ 2.2 Selection of Features
The features are carefully chosen from previous physics-based modeling research regarding adsorption energy computation. The features are categorized into three groups: adsorbate, catalyst-bulk, and surface features, as seen in Table 1. Each feature provides insights into electronic structure, bonding characteristics, and energetic properties of adsorption sites and catalysts. For instance, local electronegativity is determined as the geometric mean of the Pauling electronegativity of metal atoms within the first neighboring shell, offering insight into the local reactivity of catalyst surfaces19 as follows:
@@
where $\chi_{j}^{0}$ is the electronegativity of atom j and N is the total number of atoms within the first neighboring shell including the adsorption site i. While, site type represented numerical classifications of different adsorption site configurations such as bridge sites, ontop sites and hollow-HPC sites.
@@
Sum of adsorbate atomic numbers $\Sigma Z_{ads}$ Tran et al.21 Sum of atomic numbers of all atoms in adsorbate molecule
Center atom coordination number $CN_{center}$ Calle-Vallejo et al.22 Coordination number of the central atom in adsorbate molecule
Adsorbing atom atomic number $Z_{adatom}$ Roy et al.23 Atomic number of adsorbing atom present in adsorbate
Adsorbing atom electronegativity $\chi_{adatom}$ Amos et al.24 Pauling’s electronegativity of adsorbing atom present in adsorbate
Adsorbing atom coordination number $CN_{adatom}$ Calle-Vallejo et al.22 Coordination number of adsorbing atom present in the adsorbate
Catalyst-bulk Number of adsorbate atoms $N_{ads}$ Amos et al.24 Total number of atoms present in adsorbate molecule
@@
Formation energy $\Delta H_{f}$ Ha et al.27 Formation energy of catalyst
Mean atomic number of bulk/catalyst $\bar{z}_{bulk}$ Tran et al.21 Mean atomic number of the catalyst atoms
Fermi Energy $E_{F}$ Ishioka et al.20 Fermi Energy of catalyst
Surface Catalyst electronegativity $\chi_{cat}$ Ishioka et al.20 Average electronegativity of the interacting surface atoms of catalyst
Effective coordination number $CN_{cat}$ Calle-Vallejo et al.22 Average coordination number of the interacting surface atoms of catalyst
Site type $S$ Roy et al.23 Type of adsorption site based on interaction
Miller index $hkl$ Chen et al.28 Miller index of catalyst
Adsorbate-catalyst bond length $d_{ac}$ Ha et al.27 Bond length of adsorbing atom and slab atom
@@ 2.3 Post-hoc XAI analysis
The dataset, comprising 24053 adsorbate-catalyst systems, is divided into an 85:15 ratio for training and testing purposes, with input features and the adsorption energy label provided to the machine learning models. Details about the dataset and data selection are elaborated in the Methods section. The performance outcomes of the models are presented in Table 2. The outcomes demonstrated by all models exhibit comparable performance with MAE values ranging from 0.398 to 0.617. Notably, Adaboost regression stands out as the best performer, demonstrating the lowest MAE value. Therefore, it is used as the reference model for SHAP analysis in the post-hoc analysis. This highlights an important point: post-hoc analysis heavily depends on model performance. In other words, the analysis reflects how the model interprets the features, so advancing the model is essential to gain deeper insights.
Table 2: Model Performance Comparison Based on MAE (Mean Absolute Error) in eV
@@
In the post-hoc analysis, the primary aspect of model explainability is carried out through SHAP analysis 31. Before conducting the SHAP analysis, a correlation matrix (Figure 2(b)) was generated to assess the relationships between input features. The matrix reveals several key relationships, such as a strong positive correlation between $X_{\text{ads}}$ and $X_{\text{cat}}$, indicating potential redundancy. Conversely, features like band gap $E_{g}$ and effective coordination number $CN_{\text{cat}}$ exhibit weak correlations, suggesting independent contributions to the model. Additionally, properties like $\Delta H_{f}$ (enthalpy of formation) and $\rho_{\text{cat}}$ (density of the catalyst) show moderate correlations with other features, implying they may play a multifaceted role in adsorption energy predictions. This analysis helps in eliminating highly correlated features to reduce redundancy and improve model efficiency.
@@
The SHAP library has been widely used in various domains, including materials science, to interpret black box machine learning models 18. This technique computes Shapley values for all features, thereby elucidating each feature’s contribution to the prediction of adsorption energy by assessing the impact of each feature across different combinations of feature values.
@@
Figure 2: Performance Evaluation and Feature Correlation: (a) parity plot illustrating the performance of Adaboost Regression with Base estimator as Random Forest Regressor (Best model). MAE values are calculated for systems with H, O and C1 group. (b) Correlation matrix depicting the relationships between input features. High values of correlation coefficient have been found amongst a few features such as between Local electronegativity and Catalyst electronegativity, Site type and coordination number of adsorbate molecule.
Figure 3(a) and Figure 3(b) display the feature importance using a radar chart and a summary plot, respectively. From the feature importance, the top five features (higher Shapley values) are adsorbate electronegativity ($\chi_{\text{ads}}$), coordination number of the adsorbing atom ($CN_{\text{adatom}}$), electronegativity of the adsorbing atom ($\chi_{\text{adatom}}$), catalyst electronegativity ($\chi_{\text{cat}}$), and the number of atoms in the adsorbate ($N_{\text{ads}}$), with Shapley values of 0.46, 0.43, 0.25, 0.19, and 0.19, respectively. Electronegativity stands out as the most significant feature among the catalyst-related factors, which is consistent with findings reported in the literature (as referenced in Table 1). Three of the top five features—electronegativity of the adsorbate, adsorbing atom, and catalyst surface atoms—underscore the pivotal role of electronegativity in determining adsorption behavior on catalytic surfaces.
Conversely, a few features exhibit lower importance. For instance, Formation energy $\Delta H_{f}$, which plays a crucial role in single-atom catalysts as seen in Ha et al.27, demonstrates limited relevance in our case, which includes heterogeneous catalysts. This may be due to its nature as a bulk property, making it less effective in predicting the interactions between adsorbates and catalytic surfaces in heterogeneous systems, as reflected in its low Shapely value (+0.06). Similarly, the band gap $E_{g}$, a bulk electronic property, holds minimal predictive power for surface interactions, with a Shapely value of +0.06.
Moreover, there appears to be a stronger influence of adsorbate properties compared to catalyst properties in determining adsorption behavior. This may suggest that adsorption energy is more dependent on the type of adsorbate, possibly due to the limited variety of adsorbates, with only three categories (H, O & C1) represented in the dataset, compared to the broader range of catalysts. This imbalance in the number of adsorbates versus catalysts could influence the observed feature importance, highlighting the significant role of data distribution in shaping the outcomes of the feature importance analysis.
@@
Previous studies have shown varying relationships between catalyst electronegativity and adsorption energy. Gao, Wang, et al. proposed an inverse relationship specifically for transition metal catalysts, where higher electronegativity leads to lower adsorption energy 32. Conversely, Trasatti’s research suggests a direct proportionality between metal catalyst electronegativity and adsorption energy 33.
Our analysis reveals a positive correlation between catalyst electronegativity ($\chi_{\text{cat}}$) and adsorption energy. This result is evident in the beeswarm plot analysis shown in Figure 3(c), where significant clustering of data points and feature values for catalyst electronegativity can be observed for positive Shapley values. Unlike the findings by Gao and Wang, which show an inverse relationship between catalyst electronegativity and adsorption energy for transition metal catalysts, our dataset includes a broader range of adsorbates—hydrogen, oxygen, and carbon-based—paired with metal as well as non-metal catalysts. This diversity may account for the differences observed. This underscores the intricate nature of catalyst-adsorbate interactions, where various factors such as the composition of the dataset and the choice of adsorbates and catalysts play crucial roles in determining the observed outcomes. Additionally, a similar trend is observed for adsorbate electronegativity ($\chi_{\text{ads}}$) and local electronegativity ($\chi_{\text{local}}$), supporting our findings and emphasizing the complexity of catalyst-adsorbate interactions.
@@
In adsorption studies, surface characteristics are typically the primary focus, rather than bulk properties. However, for this dataset, the electronegativity of the catalyst bulk may still provide valuable insights into surface properties. This inference is supported by Figure 4, which explores feature-feature correlations and reveals two notable findings. Firstly, Figure 4(a) demonstrates a positive correlation between mean catalyst electronegativity ($\chi_{cat}$) and local electronegativity ($\chi_{local}$). This relationship reflects the influence of catalyst atoms within the first neighboring shell at the adsorption site. It logically follows that the electronegativity of a subset of surface atoms contributes to the overall mean electronegativity of the catalyst lattice. Additionally, a similar correlation is observed in the correlation matrix depicted in Figure 2(b), where local electronegativity and catalyst electronegativity exhibit a high correlation coefficient. This suggests a notable association between surface and bulk electronegativity in this dataset. If a strong correlation between bulk and surface characteristics is established for the target dataset, bulk characteristics could serve as proxies for surface properties, potentially offering greater flexibility in data acquisition, particularly for experimental studies.
The increase in the number of adsorbate atoms correlates with a decrease in overall adsorbate electronegativity, as depicted in Figure 4(b). This trend is evident from the transition in color scheme from orange (indicating high adsorbate electronegativity) to purple (indicating lower adsorbate electronegativity). The limited variety of adsorbate types in our dataset likely amplifies the role of adsorbates as key determinants in the adsorption process. The decrease in electronegativity could be attributed to the formation of chemical bonds between the adsorbate atoms. As more adsorbate atoms come into proximity, bonding interactions occur, causing a redistribution of electron density within the adsorbate system and thereby lowering the overall electronegativity 34.Adsorbate bonding and composition are key to catalyst interactions, directly affecting adsorption efficiency. This insight can guide the construction of more robust and targeted datasets for adsorption studies.